# Probing cluster magnetism with embedded and background radio sources in Planck clusters

### v1

@@
22institutetext: Leiden Observatory, Leiden University, PO Box 9513, NL-2300 RA Leiden, The Netherlands 33institutetext: Minnesota Institute for Astrophysics, University of Minnesota, 116 Church St SE, Minneapolis, MN 55455, USA 44institutetext: Center for Astrophysics — Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA 55institutetext: Department of Liberal Arts and Sciences, Berklee College of Music, 7 Haviland Street, Boston, MA 02215, USA 66institutetext: DIFA - Università di Bologna, via Gobetti 93/2, I-40129 Bologna, Italy 77institutetext: INAF - IRA, Via Gobetti 101, I-40129 Bologna, Italy; IRA - INAF, via P. Gobetti 101, I-40129 Bologna, Italy 88institutetext: Naval Research Laboratory, 4555 Overlook Avenue SW, Code 7213, Washington, DC 20375, USA 99institutetext: Institute for Astronomy, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, UK
@@
E. Osinga 1122    R. J. van Weeren 22    L. Rudnick 33    F. Andrade-Santos 4455    A. Bonafede 6677    T. Clarke 88    K. Duncan 99    S. Giacintucci 88    H. J. A. Röttgering 22
@@
Magnetic fields remain an enigmatic part of the content of galaxy clusters. Faraday rotation and depolarisation of extragalactic radio sources are useful probes, but the limited availability of polarised radio sources necessitates the stacking of clusters to study average magnetic field profiles and correlation scales. We recently presented a Karl G. Jansky Very Large Array survey of the 124 most massive Planck clusters at low redshift ($z<0.35$), where a clear depolarisation trend with the cluster impact parameter was found, with sources at smaller projected distances to the cluster centre showing more depolarisation. In this study, we combine the depolarisation information with the observed rotation measure (RM) and present an investigation into the average magnetic field properties of the sample, using both background sources and sources embedded in clusters. We observe a significant increase in the RM scatter, $\sigma_{\mathrm{RRM}}$, closer to the cluster centres. Averaging all 124 clusters, we find a scatter within $R_{\mathrm{500}}$ of $\sigma_{\mathrm{RRM}}=209\pm 37$ rad m-2, with background sources and cluster members showing similar values ($200\pm 33$ and $219\pm 66$ rad m-2, respectively). In the simple assumption of a uniform magnetic field with a single fluctuation scale $\Lambda_{c}$, this translates to an average magnetic field strength of $2\,(\Lambda_{c}/10\mathrm{kpc})^{-0.5}\,\mu$G. The profile of $\sigma_{\mathrm{RRM}}$ as a function of projected radius is inconsistent with a model that has a simple scaling $B\propto n_{e}^{\eta}$, with an observed deficit near the centre of clusters possibly caused by the fact that the highest RM sources near the centre of clusters are depolarised. Combining depolarisation and RM in a full forward model, we find that the magnetic field power spectrum roughly agrees with the Kolmogorov value, but that none of the Gaussian random field models can fully explain the observed relatively flat profiles. This implies that more sophisticated models of cluster magnetic fields in a cosmological context are needed.
@@ 1 Introduction
The most promising tool to study magnetic fields is radio polarisation observations of Faraday rotation and depolarisation. A magnetised plasma such as the ICM causes a wavelength-dependent rotation of the polarisation angle. In general, the Faraday depth of a source is defined as (Burn, 1966; Brentjens & de Bruyn, 2005)
@@
where $n_{e}$ is the electron density in parts per cm-3, B is the magnetic field in $\mu$Gauss and dr the infinitesimal path length increment along the line of sight (LOS) in kpc, and we define $\phi(\textbf{r})>0$ for the magnetic field pointing towards the observer. In the simple case of just one radio-emitting source along the line of sight, the Faraday depth is equal to the rotation measure (RM). With a combination of radio observations of RMs and X-ray observations of electron densities, it is thus possible to study the magnetic field properties of galaxy clusters.
Such studies are best done at low redshifts to maximise the cluster angular size because polarised radio sources are relatively rare (e.g. Rudnick & Owen, 2014). The most detailed analyses have been of the Coma Cluster (Bonafede et al., 2010) and Abell 2345 (Stuardi et al., 2021), where seven polarised radio sources were detected per cluster. The Coma Cluster magnetic field was found to agree with a Kolmogorov power spectrum with a central strength of 5 $\mu$G and a scaling of magnetic field energy density linearly proportional to the thermal gas density ($B^{2}\propto n_{e}$). The central magnetic field strength in Abell 2345 was found to be similar to the Coma Cluster, but with a magnetic field energy density that scales super-linearly instead (i.e. $B^{2}\propto n_{e}^{2}$). Several other low-redshift clusters have been analysed in polarisation (Murgia et al., 2004; Govoni et al., 2006; Guidetti et al., 2008; Govoni et al., 2010; Vacca et al., 2012; Govoni et al., 2017), with typically less than five polarised radio galaxies per study, resulting in large uncertainties on the magnetic field estimates (see e.g. Johnson et al., 2020, for a detailed discussion).
@@
We recently published the largest homogeneous stacking experiment using Karl G. Jansky Very Large Array (VLA) observations of 124 galaxy clusters (Osinga et al., 2022) selected from the Planck 2nd Sunyaev-Zeldovich Source Catalog (PSZ2) (Planck Collaboration et al., 2016). This study presented the first clear depolarisation trend tracing the radial profile of cluster magnetic fields using over 600 polarised radio sources. While depolarisation traces the smaller scale structure of the magnetic fields (i.e. few kpc), the larger scale structure can be inferred from the Faraday rotation of the radio sources. In this paper, we add the information from the Faraday rotation of the same sample of sources to study the large-scale properties of the magnetic fields in galaxy clusters. By jointly fitting both depolarisation and Faraday rotation, we aim to constrain the average magnetic field strength, scaling with density, and power spectrum. Cosmological calculations are performed assuming a flat $\Lambda$CDM model with $H_{0}=70$ kms-1Mpc-1, $\Omega_{m}=0.3$ and $\Omega_{\Lambda}=0.7$ and the $R_{\mathrm{500}}$ radius is the radius enclosing an overdensity of $500$ at the cluster redshift.
@@ 2 Chandra-Planck ESZ sample
Each of the 124 clusters was observed for $\sim 40$ min in the VLA L-band (1–2 GHz), resulting in typically 20–30 $\mu\mathrm{Jy\,beam}^{-1}$ noise levels at a resolution of 6–7′′  after data reduction. Polarised sources were identified using RM-synthesis (Brentjens & de Bruyn, 2005) and matched to total intensity components and optical counterparts. Optical counterparts were used to determine cluster membership, with spectroscopic redshifts where available and photometric redshifts otherwise. In total 6,807 and 819 source components were detected in total and polarised intensity respectively. Throughout this paper we will simply refer to these components as ‘sources’. We have fit the following model to the polarised intensity as a function of wavelength $\lambda$, which accounts for rotation and depolarisation (see Sokoloff et al., 1998, for details),
@@
where $p_{0}$ denotes the intrinsic polarisation and $\Sigma^{2}_{\mathrm{RM}}$ the inferred variance of the RM distribution within the observing beam which models the depolarisation as a function of wavelength [^2] We note that Osinga et al. (2022) defined this as $\sigma^{2}_{\mathrm{RM}}$, but we use a different nomenclature here to avoid confusion with the scatter in RM between sources.. $\chi_{0}$ is the intrinsic polarisation angle, and $I$ denotes the total intensity model, which was assumed to be a simple power-law of the form $I(\nu)=I_{0}\nu^{\alpha}$ where $\nu$ is the frequency and $\alpha$ the spectral index.
We have improved the Monte Carlo Markov chain (MCMC) fitter used by Osinga et al. (2022) to now properly take into account the circular nature of $\chi_{0}$ during the fitting. Osinga et al. (2022) used a uniform prior on $\chi_{0}$, $\sim\mathcal{U}(0,\pi)$, which would cause the sampler in some cases to become trapped around the boundary values. We removed this prior on $\chi_{0}$ and fold the chain back into the range $[0,\pi)$ after the sampling is completed. We also calculate the mean and spread using circular statistics where we take into account the fact that the angles are distributed on the half-circle $\in$ $[0,\pi)$. In this way, the mean denotes the angle of the average vector on the unit circle, and the standard deviation is the spread in angles around the average vector. This agrees with the definition of the simple arithmetic mean and standard deviation when the angles are distributed away from the edges of the domain.
Table 1: Number counts of polarised sources and unique clusters in which they were detected for different subsets of the catalogue in Appendix C. Only clusters with X-ray data were forward-modelled (Sec. 4.4 & 4.5).
@@
This paper is accompanied by an updated table of polarised components, shown in Appendix C. We note that this update mainly corrects the quoted mean and uncertainty of the intrinsic polarisation angle $\chi_{0}$ and most sources have similar best-fit RM and depolarisation parameters. This thus does not significantly impact the results. Finally, we used the same criteria for identifying bad fits as Osinga et al. (2022). All sources with a best-fit $\chi^{2}$ value that is $>5\sigma$ away from the theoretical distribution, and sources with low signal-to-noise polarised emission resulting in artificially large values of $\Sigma_{\mathrm{RM}}^{2}$ were flagged. This led to 195 bad fits among the 819 polarised sources identified in 124 galaxy clusters. After excluding three clusters (G121.11+57.01, G115.71+17.52, G139.59+24.18) due to a lack of accurate redshift estimates, a total of 610 sources across 117 clusters were available for subsequent analysis. Table 1 provides an overview of the number counts for various sub-samples and the unique clusters in which these sources were found, aiding in understanding the selection function of the defined subsets in the following analysis.
@@ 3.1 Determining RM scatter
  • $\delta\sigma_{\mathrm{RM}}$ for uncertainty on the scatter in RM.
@@
$\begin{split}\mathrm{RM}_{\mathrm{obs^{\prime}}}&=\frac{\mathrm{RM}_{\mathrm{src}}}{(1+z_{\mathrm{intr}})^{2}}+\mathrm{RM}_{\mathrm{IGM}}+\frac{\mathrm{RM}_{\mathrm{cluster}}}{(1+z_{\mathrm{cluster}})^{2}}\\ &\quad+\mathrm{RM}_{\mathrm{gal}}+\mathrm{RM}_{\mathrm{ion}}\end{split}$ (3)
The scatter in RM between sources, or RM variance, analogously is the sum of the variance of these individual terms, as they are independent contributions. The Galactic foreground can be largely removed by correcting the RMs for the Galactic contribution, which has been mapped most recently by Hutschenreuter et al. (2022). The measurement noise is known from the estimated uncertainties in RM from the MCMC spectral fitting. The most difficult components to isolate are the variance caused by the IGM and the intrinsic RM variance of radio sources. Studies using low RM variance sources (i.e. away from clusters and groups) that attempt to isolate the extragalactic from the Galactic and intrinsic scatter find that the extragalactic RM (rms) scatter is between $\sim$ 1 and 10 rad m-2 (Oppermann et al., 2015; Vernstrom et al., 2019; O’Sullivan et al., 2020; Pomakov et al., 2022).
@@
$\mathrm{RRM}=\frac{\mathrm{RM}_{\mathrm{cluster}}}{(1+z_{\mathrm{cluster}})^{2}}+\delta\mathrm{RRM}+\mathrm{RM}_{\mathrm{extr}}$ (4)
where we defined the last term as the extrinsic RM, which captures the RM contribution that is not from the cluster or the measurement uncertainties. The extrinsic RM consists of components from the IGM, the radio source itself, and an ionospheric component, which are difficult to separate. However, we can estimate the scatter caused by $\mathrm{RM}_{\mathrm{extr}}$ through sources located far away from clusters. At large radii ($r>3R_{\mathrm{500}}$), the scatter is expected to be unrelated to the ICM, and caused mainly by the extrinsic scatter between sources and noise scatter due to measurement uncertainties. We find that the standard deviation of the observed RMs corrected for the Galactic contribution is equal to 13 $\pm$ 2 rad m-2 (see Fig. 18), with the measurement uncertainties accounting for 9 rad m-2. This gives an estimate of $\sigma_{\mathrm{RM,extr}}=9$ rad m-2, which is of the same order as the observed dispersion of extragalactic RMs (e.g. Schnitzeler, 2010; Vernstrom et al., 2019).

### v2

@@
22institutetext: Leiden Observatory, Leiden University, PO Box 9513, NL-2300 RA Leiden, The Netherlands 33institutetext: Minnesota Institute for Astrophysics, University of Minnesota, 116 Church St SE, Minneapolis, MN 55455, USA 44institutetext: Center for Astrophysics — Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA 55institutetext: Department of Liberal Arts and Sciences, Berklee College of Music, 7 Haviland Street, Boston, MA 02215, USA 66institutetext: Clay Center Observatory, Dexter Southfield, 20 Newton Street, Brookline, MA 02445, USA 77institutetext: DIFA - Università di Bologna, via Gobetti 93/2, I-40129 Bologna, Italy 88institutetext: INAF - IRA, Via Gobetti 101, I-40129 Bologna, Italy; IRA - INAF, via P. Gobetti 101, I-40129 Bologna, Italy 99institutetext: Naval Research Laboratory, 4555 Overlook Avenue SW, Code 7213, Washington, DC 20375, USA 1010institutetext: Institute for Astronomy, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, UK
@@
E. Osinga 1122    R. J. van Weeren 22    L. Rudnick 33    F. Andrade-Santos 445566    A. Bonafede 7788    T. Clarke 99    K. Duncan 1010    S. Giacintucci 99    H. J. A. Röttgering 22
@@
Magnetic fields remain an elusive part of the content of galaxy clusters. Faraday rotation and depolarisation of extragalactic radio sources are useful probes, but the limited availability of polarised radio sources necessitates the stacking of clusters to study average magnetic field properties. We recently presented a Karl G. Jansky Very Large Array survey of the 124 most massive Planck clusters at low redshift ($z<0.35$), finding a clear depolarisation trend with the cluster impact parameter, with sources at smaller projected distances to the cluster centre showing more depolarisation. In this study, we combine the depolarisation information with the observed rotation measure (RM) and present an investigation into the average magnetic field properties of the sample, using both background sources and sources embedded in clusters. We observe a significant increase in the RM scatter, $\sigma_{\mathrm{RRM}}$, closer to the cluster centres. Averaging all 124 clusters, we find a scatter within $R_{\mathrm{500}}$ of $\sigma_{\mathrm{RRM}}=209\pm 37$ rad m-2, with background sources and cluster members showing similar values ($200\pm 33$ and $219\pm 66$ rad m-2, respectively). In the simple assumption of a uniform amplitude magnetic field with a single fluctuation scale $\Lambda_{c}$, this translates to an average magnetic field strength of $2\,(\Lambda_{c}/10\mathrm{kpc})^{-0.5}\,\mu$G. The profile of $\sigma_{\mathrm{RRM}}$ as a function of projected radius is inconsistent with a model that has a simple scaling $B\propto n_{e}^{\eta}$, with an observed deficit near the centre of clusters possibly caused by the fact that the highest RM sources near the centre of clusters are depolarised. Combining depolarisation and RM in a full forward model, we find that the magnetic field power spectrum roughly agrees with the Kolmogorov value, but that none of the Gaussian random field models can fully explain the observed relatively flat profiles. This implies that more sophisticated models of cluster magnetic fields in a cosmological context are needed.
@@ 1 Introduction
The most promising tool to study magnetic fields is radio polarisation observations. A magnetised plasma such as the ICM causes a wavelength-dependent rotation of the polarisation angle (Faraday rotation) and depolarisation. In general, the Faraday depth of a source is defined as (Burn, 1966; Brentjens & de Bruyn, 2005)
@@
where $n_{e}$ is the electron density in parts per cm-3, B is the magnetic field in $\mu$Gauss and dr the infinitesimal path length increment along the line of sight (LOS) in kpc, and we define $\phi(\textbf{r})>0$ for the magnetic field pointing towards the observer. In the simple case of just one radio-emitting source along the line of sight, the Faraday depth is equal to the rotation measure (RM). With a combination of RMs from radio observations and electron densities from X-ray observations, it is thus possible to study the magnetic field properties of galaxy clusters.
Such studies are best done at low redshifts to maximise the cluster angular size because polarised radio sources are relatively rare (e.g. Rudnick & Owen, 2014). The most detailed analyses have been of the Coma Cluster (Bonafede et al., 2010) and Abell 2345 (Stuardi et al., 2021), where seven polarised radio sources were detected per cluster. The Coma Cluster magnetic field was found to agree with a Kolmogorov power spectrum with a central strength of 5 $\mu$G and a scaling of magnetic field energy density linearly proportional to the thermal gas density ($B^{2}\propto n_{e}$). The central magnetic field strength in Abell 2345 was found to be similar to the Coma Cluster, but with a magnetic field energy density that scales more steeply instead (i.e. $B^{2}\propto n_{e}^{2}$). Several other low-redshift clusters have been analysed in polarisation (Murgia et al., 2004; Govoni et al., 2006; Guidetti et al., 2008; Govoni et al., 2010; Vacca et al., 2012; Govoni et al., 2017), with typically less than five polarised radio galaxies per study, resulting in large uncertainties on the magnetic field estimates (see e.g. Johnson et al., 2020, for a detailed discussion).
@@
We recently published the largest homogeneous stacking experiment using Karl G. Jansky Very Large Array (VLA) observations of 124 galaxy clusters (Osinga et al., 2022) selected from the Planck 2nd Sunyaev-Zeldovich Source Catalog (PSZ2) (Planck Collaboration et al., 2016). This study presented the first clear depolarisation trend tracing the radial profile of cluster magnetic fields using over 600 polarised radio sources. While depolarisation traces the smaller scale structure of the magnetic fields on the scale of the resolving beam (usually kpc), the larger scale structure (i.e. along the line of sight, usually Mpc) can be inferred from the Faraday rotation of the radio sources. In this paper, we add the information from the Faraday rotation of the same sample of sources to study the large-scale properties of the magnetic fields in galaxy clusters. By jointly fitting both depolarisation and Faraday rotation, we aim to constrain the average magnetic field strength, scaling with density, and power spectrum. Cosmological calculations are performed assuming a flat $\Lambda$CDM model with $H_{0}=70$ kms-1Mpc-1, $\Omega_{m}=0.3$ and $\Omega_{\Lambda}=0.7$ and the $R_{\mathrm{500}}$ radius is the radius enclosing an overdensity of $500$ at the cluster redshift.
@@ 2 Chandra-Planck ESZ sample
Each of the 124 clusters was observed for $\sim 40$ min in the VLA L-band (1–2 GHz), resulting in typically 20–30 $\mu\mathrm{Jy\,beam}^{-1}$ noise levels at a resolution of 6–7′′  after data reduction. Polarised sources were identified using RM-synthesis (Brentjens & de Bruyn, 2005) and matched to total intensity components and optical counterparts. Optical counterparts were used to determine cluster membership, with spectroscopic redshifts where available and photometric redshifts otherwise. The full details on redshift determination are given in Osinga et al. (2022), where Eq. 25-26 were used to divide sources into background and cluster members. In total 6,807 and 819 source components were detected in total and polarised intensity respectively. Throughout this paper we will simply refer to these components as ‘sources’. We have fit the following model to the polarised intensity as a function of wavelength $\lambda$, which accounts for rotation and depolarisation
@@
where $p_{0}$ denotes the intrinsic polarisation and $\Sigma^{2}_{\mathrm{RM}}$ the inferred variance of the RM distribution within the observing beam which models the depolarisation as a function of wavelength [^2] We note that Osinga et al. (2022) defined this as $\sigma^{2}_{\mathrm{RM}}$, but we use a different nomenclature here to avoid confusion with the scatter in RM between sources.. $\chi_{0}$ is the intrinsic polarisation angle, and $I$ denotes the total intensity model, which was assumed to be a simple power-law of the form $I(\nu)=I_{0}\nu^{\alpha}$ where $\nu$ is the frequency and $\alpha$ the spectral index. Equation 2 assumes that (see Sokoloff et al., 1998, for details):
  1. 1.
    The dominant mechanism causing Faraday rotation is an external Faraday screen (i.e. a non-emitting foreground magnetised plasma, such as the ICM).
  2. 2.
    The foreground screen magnetic field changes on scales (at least) smaller than the observing beam, which causes Faraday dispersion inside the beam, resulting in depolarisation.
  3. 3.
    The magnetic field fluctuates as a Gaussian random field, implying its scatter can be completely captured in the parameter $\Sigma_{\mathrm{RM}}$.
It has been shown (Osinga et al., 2022) that this model accurately fits 75% of the polarised radio sources detected in this sample, and we discard sources that do not show a proper fit (see below).
We have improved the Monte Carlo Markov chain (MCMC) fitter used by Osinga et al. (2022) to now properly take into account the circular nature of $\chi_{0}$ during the fitting. Osinga et al. (2022) used a uniform prior on $\chi_{0}$, $\sim\mathcal{U}(0,\pi)$, which would cause the sampler in some cases to become trapped around the boundary values. We removed this prior on $\chi_{0}$ and fold the chain back into the range $[0,\pi)$ after the sampling is completed. We also calculate the mean and spread using circular statistics where we take into account the fact that the angles are distributed on the half-circle $\in$ $[0,\pi)$. In this way, the mean denotes the angle of the average vector on the unit circle, and the standard deviation is the spread in angles around the average vector. This agrees with the definition of the simple arithmetic mean and standard deviation when the angles are distributed away from the edges of the domain (see e.g. Mardia & Jupp, 2009, for more details on circular statistics).
Table 1: Number counts of polarised sources and unique clusters in which they were detected for different subsets of the catalogue in Appendix C.
@@
 [^3] Only clusters with X-ray data were forward-modelled (Sec. 4.4 & 4.5).
This paper is accompanied by an updated table of polarised components, shown in Appendix C. We note that this update mainly corrects the quoted mean and uncertainty of the intrinsic polarisation angle $\chi_{0}$ and most sources have similar best-fit RM ($\phi$) and depolarisation ($\Sigma_{\mathrm{RM}}$) parameters to within a few per cent. This thus does not significantly impact the results. Finally, we used the same criteria for identifying bad fits as Osinga et al. (2022). All sources with a best-fit $\chi^{2}$ value that is $>5\sigma$ away from the theoretical distribution, and sources with low signal-to-noise polarised emission resulting in artificially large values of $\Sigma_{\mathrm{RM}}^{2}$ were flagged. This led to 195 bad fits among the 819 polarised sources identified in 124 galaxy clusters. After excluding three clusters (G121.11+57.01, G115.71+17.52, G139.59+24.18) due to a lack of accurate redshift estimates, a total of 610 sources across 117 clusters were available for subsequent analysis. Table 1 provides an overview of the number counts for various sub-samples and the unique clusters in which these sources were found, aiding in understanding the selection function of the defined subsets in the following analysis.
@@ 3.1 Determining RM scatter
  • $\delta\sigma_{\mathrm{RM}}$ for uncertainty on the scatter in RM ($\sigma_{\mathrm{RM}}$).
@@
$\begin{split}\mathrm{RM}_{\mathrm{obs^{\prime}}}&=\frac{\mathrm{RM}_{\mathrm{src}}}{(1+z_{\mathrm{intr}})^{2}}+\mathrm{RM}_{\mathrm{IGM}}+\frac{\mathrm{RM}_{\mathrm{cluster}}}{(1+z_{\mathrm{cluster}})^{2}}\\ &\quad+\mathrm{RM}_{\mathrm{gal}}+\mathrm{RM}_{\mathrm{ion}},\end{split}$ (3)
where $\mathrm{RM}_{\mathrm{IGM}}$ is the integrated contribution of the IGM, which itself has a redshift dependence (c.f. Amaral et al., 2021, Eq. 2) that is however not relevant for this work as we regard it as a single nuisance parameter.
The scatter in RM between sources, or RM variance, analogously is the sum of the variance of these individual terms, as they are independent contributions. The measurement noise is known from the estimated uncertainties in RM from the MCMC spectral fitting (we use the $16$th and $84$th percentile of the posterior, see Osinga et al., 2022, for details). The Galactic foreground can be largely removed by correcting the RMs for the Galactic contribution, which has been mapped most recently by Hutschenreuter et al. (2022). The most difficult components to isolate are the variance caused by the IGM and the intrinsic RM variance of radio sources. Studies using low RM variance sources (i.e. away from clusters and groups) that attempt to isolate the extragalactic from the Galactic and intrinsic scatter find that the extragalactic RM (rms) scatter is between $\sim$ 1 and 10 rad m-2 (Oppermann et al., 2015; Vernstrom et al., 2019; O’Sullivan et al., 2020; Pomakov et al., 2022).
@@
$\mathrm{RRM}=\frac{\mathrm{RM}_{\mathrm{cluster}}}{(1+z_{\mathrm{cluster}})^{2}}+\mathrm{RM}_{\mathrm{extr}}$ (4)
where we defined the last term as the extrinsic RM, which captures the RM contribution that is not from the cluster or the measurement uncertainties ($\delta\mathrm{RRM}$). The extrinsic RM consists of components from the IGM, the radio source itself, and an ionospheric component, which are difficult to separate. However, we can estimate the scatter caused by $\mathrm{RM}_{\mathrm{extr}}$ through sources located far away from clusters. At large radii ($r>3R_{\mathrm{500}}$), the scatter is expected to be unrelated to the ICM, and caused mainly by the extrinsic scatter between sources and noise scatter due to measurement uncertainties. We find that the standard deviation of the observed RMs corrected for the Galactic contribution is equal to 13 $\pm$ 2 rad m-2 (see Fig. 18), with the measurement uncertainties accounting for 9 rad m-2. This gives an estimate of $\sigma_{\mathrm{RM,extr}}=9$ rad m-2, which is of the same order as the observed dispersion of extragalactic RMs (e.g. Schnitzeler, 2010; Vernstrom et al., 2019).